OpenStudy (anonymous):

and yet another question...... A car is driving northwest at v mph across a sloping plain whose height, in feet above sea level, at a point N miles north and E miles east of a city is given by h(N,E)=3000+75N+75E. (a) At what rate is the height above sea level changing with respect to distance in the direction the car is driving? rate = (b) Express the rate of change of the height of the car with respect to time in terms of v. rate =

OpenStudy (anonymous):

grad f =<75,75>, so the rate of change is the mag of that....well, i put in sqrt (75^2+75^2> and webwork said no.............

OpenStudy (yanasidlinskiy):

A.) \(dh \div dt=75\times dN \div dt +100 dE \div dt\) where, \(dN \div dt =V \cos 45\) \(dE/dt=-V \cos 45\) Find: \(dh \div dt\) Rate is the height above sea level changing with respect to distance in the direction the car is driving= rate is the height above sea level changing with respect to distance in the direction the car is driving= \(V -dh \div dt\) Solve:)

OpenStudy (anonymous):

those aren't even the numbers i need to work with....and this is something that was just copy and pasted from yahoo answers that i looked at and was incorrect.